OpenStudy (jamesreed30):

Find the exact circumference in terms of PI of a circle with the given radius. 7 feet

OpenStudy (mathmale):

Find the formula for the circumference of a circle. C=? Replace "r" (the radius) in that formula with 7 ft. End.

OpenStudy (dallascowboys88):

Yes I Agree With @mathmale

OpenStudy (jamesreed30):

So C=2 PI? Then what do I do?

OpenStudy (dallascowboys88):

@mathmale is the one to help you he's very very smart so remember to ask him for help nextime

OpenStudy (jamesreed30):

I'm new to this site how do I ask just him?

OpenStudy (jamesreed30):

@mathmale

OpenStudy (jamesreed30):

I'm new to this site how do I ask just him?

OpenStudy (dallascowboys88):

just what you just did just say, @mathmale and it will tell him you need his help.

OpenStudy (mathmale):

Your formula is partially correct, but still needs work. Please go back, find the formula for the circumference of a circle, and share it here.

OpenStudy (jamesreed30):

Ok thank you.

OpenStudy (dallascowboys88):

np thx @mathmale for you help too.

OpenStudy (mathmale):

The formula for the circumference of a circle of radius "r" is .... ?

OpenStudy (jamesreed30):

7?

OpenStudy (mathmale):

Have you looked up this formula, James? Please Google "circumference of a circle" and copy down the relevant formula.

OpenStudy (jamesreed30):

Alright

OpenStudy (mathmale):

C = ??

OpenStudy (jamesreed30):

C= 2pi r

OpenStudy (jamesreed30):

@mathmale

OpenStudy (jamesreed30):

@brooke..help00

OpenStudy (mathmale):

Good. Now substitute "7 feet" for "r" in that formula. Share your result, please.

OpenStudy (jamesreed30):

Do I divide or multiply 7 and 2? @mathmale

OpenStudy (mathmale):

No. The formula is C=2 pi r. 2 pi r means "2 times pi times radius r." Write out C = 2 pi r, and then write it again, substituting '7 feet' for r. Show your work, please.

OpenStudy (jamesreed30):

So C= 2 PI 7. Then if 2 times PI times radius 2 x 3.14?

OpenStudy (mrnood):

This is exactly the same question that oyu asked yesterday with a different value for r The answer method is the same it asks for it in 'in terms of pi' you need pi IN your answer

OpenStudy (jamesreed30):

It include PI for me, I have to write it a weird way

OpenStudy (mrnood):

no - you have put in a number for pi (3.14) - you must leave "pi" in your answer

OpenStudy (mrnood):

but you must learn from asking these questions refer to yesterdays answer and do the same thing

OpenStudy (mrnood):

3.14 is NOT exact - it is an approximation "pi" means 'the exact value of pi'

OpenStudy (jamesreed30):

It's because it says exact answer and I don't know what exact answer means

OpenStudy (mathmale):

In this case, C = 2 pi r. Better written in mathematical terms, this would be \[C = 2 \pi r\]

OpenStudy (mrnood):

the answer to yesterdays question was exactly 10 pi now do the same but let r=7

OpenStudy (mathmale):

The problem instructions say to express your answer in terms of 'pi,' which means in terms of the symbol \[\pi\]

OpenStudy (mathmale):

So, write out \[C=2 \pi r\]

OpenStudy (mathmale):

and then replace that "r" with "7 feet." DONE.

OpenStudy (jamesreed30):

So my answer would be 12

OpenStudy (mathmale):

"the answer to yesterdays question was exactly 10 pi" TRUE. Just replace that "pi" with the symbol \[" \pi "\]

OpenStudy (jamesreed30):

I don't have to write it out in an equation it gives me that I just need a number

OpenStudy (mathmale):

I disagree. You are to write out the answer in terms of the mathematical character \[\pi\]

OpenStudy (jamesreed30):

Is there a way I can take a picture and put it on here?

OpenStudy (mathmale):

The sooner you understand this "in terms of" stuff, the better. simplify the following: \[C= 2 \pi (14~feet)\]

OpenStudy (jamesreed30):

C=2 pi (7 feet)

OpenStudy (mathmale):

and you will be done. Take a screen shot and share the resulting image. Or take a photo of the screen with your cell phone, upload the image to your computer, and then upload the image to OpenStudy. Please stop arguing. Learn what "in terms of" means, and also learn what "in terms of pi" means. It does NOT, repeat, NOT, state that you are supposed to "write a number" for pi.

OpenStudy (jamesreed30):

How do I post screenshots on here

OpenStudy (mrnood):

the only way you can write pi exactly is π you answer must include that symbol π (or "pi" will suffice in this board) (the alt codes don't seem to work here ALT227 should produce π)

OpenStudy (mrnood):

We don't need a screenshot - we understand the question You need to understand what the answer looks like yesterdays correct answer was 2π*5 = 10π "10π" is the correct exact answer now do exactly that but use 7 instead of 5 for the radius

OpenStudy (mathmale):

The circumference of a circle of radius 7 feet is ... (you fill in the answer).

OpenStudy (jamesreed30):

So 12

OpenStudy (jamesreed30):

No 14

OpenStudy (mrnood):

yesterdays correct answer was 2π*5 = 10π "10π" is the correct exact answer now do exactly that but use 7 instead of 5 for the radius Note there is a π in the answer there MUST be a π in YOUR answer π IS a number

OpenStudy (mrnood):

2π*5 = 10π change the 5 to a 7 and rewrite the answer

OpenStudy (jamesreed30):

I get there needs to be the symbol in there I have that I just don't know what the answer is like do I do 7 times 14

OpenStudy (jamesreed30):

7 times 2 my bad

OpenStudy (mrnood):

yes

OpenStudy (mrnood):

I repeat my last post do EXACTLY this and you will have your answer....> 2π*5 = 10π change the 5 to a 7 and rewrite the answer

OpenStudy (jamesreed30):

Okay so my next question is the same just with a radius of 3 1/2 so I would just do 3 1/2 times 2 get that number then put my symbol?

OpenStudy (mathmale):

Please take a chance: Review our past discussion and then DO the work.

OpenStudy (mathmale):

So much depends upon your learning something well enough to be able to apply it to a new situation. That's what you have to do here.

OpenStudy (mrnood):

Please post your answer to the question we are helping with C=2πr is the same as C=2rπ so the answer is yes it is always 2 times radius times π

OpenStudy (jamesreed30):

So then I will take 14 and do 14x 3.14?

OpenStudy (jamesreed30):

Or just leave it as 14

OpenStudy (mathmale):

Kindly DO the work and SHARE it here. Why are you using 3.14 again? STICK with "pi" or \[\pi\]

OpenStudy (mathmale):

Thre is NO reason to use 3.14 as an approx. value for pi, not here, anyway.

OpenStudy (jamesreed30):

Ight bruh. So my answer is 14 with re symbol correct?

OpenStudy (jamesreed30):

The*

OpenStudy (mrnood):

do you understand that π represents an actual number that CANNOT be written any other way exactly? the reason for NOT using 3.14 is because pi is not EQUAL to 3.14

OpenStudy (mathmale):

What was the radius? What are the correct units of measurement? Please type out the entire formula: C = 2 pi r and then make the appropriate substitution for the r value (not the pi value)./

OpenStudy (jamesreed30):

Yea I get that part man

OpenStudy (jamesreed30):

C=2 PI 7

OpenStudy (mathmale):

That's better! But what does the '7' represent? MUST show a unit of measurement.

OpenStudy (jamesreed30):

Radius

OpenStudy (mathmale):

C= 2 pi ( ? ) You MUST provide a numerical value (with units of measurement included) for the radius.

OpenStudy (jamesreed30):

7 PI ft.

OpenStudy (mathmale):

What happened to the label, "C?" What happened to the multiplier, "2?"

OpenStudy (mathmale):

C = 2 pi (7 feet) Simplify this, please.

OpenStudy (mathmale):

(Multiply '7 feet' by 2.)

OpenStudy (jamesreed30):

So 14